Lakes that have been acidified by acid rain can be neutralized by the addition of limestone. How much limestone in kg is required to completely neutralize a 4.3billion litre lake with ph of 5.5


1
Expert's answer
2021-03-29T06:10:03-0400

From the given pH, we calculate the concentration of H+:

   "[H^+] = 10^{-pH} = 10^{-5.5}"


We then use the volume to solve for the number of moles of H+:

   moles "H^+ = 10^{-5.5}M \u00d7(4.3\u00d710^9)L = 13598 moles"


From the balanced equation of the neutralization of hydrogen ion by limestone written as

   CaCO3(s) + 2H+(aq) → Ca2+(aq) + H2CO3(aq)


we use the mole ratio of limestone CaCO3 and H+ from their coefficients, which is 1 mole of CaCO3 is to react with 2 moles of H+, to compute for the mass of the limestone:

   

moles CaCO3 "= \frac{13598}{2}=6799g"

  

Since molar mass of CaCO3= 100


 mass CaCO3 "=\frac{6799×100}{1000}= 679.9 Kg"